𝑦^1/2 𝑑𝑦/𝑑π‘₯+𝑦3/2=1; 𝑦(0)=4




1
Expert's answer
2021-02-24T12:40:12-0500

Given: "y^{\frac{1}{2}}\frac{dy}{dx}+y^{\frac{3}{2}}=1"

Require to solve the given differential equation.

Now "y^{\frac{1}{2}}\frac{dy}{dx}+y^{\frac{3}{2}}=1\Rightarrow y^{\frac{1}{2}}\frac{dy}{dx}=1-y^{\frac{3}{2}}"

"\Rightarrow \frac{dy}{dx}=\frac{1-y^{\frac{3}{2}}}{y^{\frac{1}{2}}}"

By using separation of variables, we get

"\frac{y^{1/2}}{1-y^{3/2}}dy=dx"

Intetrating on both sides, we get

"\int \frac{y^{1/2}}{1-y^{3/2}}dy=\int dx"

To evaluate the above integral, let us take the substitution

"u=1-y^{3/2}\Rightarrow du=[0-\frac{3}{2}y^{1/2}]dy\Rightarrow du=-\frac{3}{2}y^{1/2}dy"

"\Rightarrow y^{1/2}dy=-\frac{2}{3}du"

Then the integral becomes

"\int \frac{-\frac{2}{3}du}{u}=\int dx"

"-\frac{2}{3}ln\left | u \right |=x+c"

Substituting "u=1-y^{3\/2}" , we get

"-\frac{2}{3}ln\left | 1-y^{3/2} \right |=x+c"

Use the initial condition "y(0)=4" to find the value of c.

"y(0)=4\Rightarrow -\frac{2}{3}ln\left | 1-4^{3/2} \right |=0+c\Rightarrow c=-\frac{2}{3}ln(7)"

Then, solution to the given differential equation is

"-\frac{2}{3}ln\left | 1-y^{3/2} \right |=x-\frac{2}{3}ln(7)"

Therefore general solution is

"\frac{2}{3}ln\left | \frac{1-y^{3/2}}{7} \right |=-x"